NottheMonga2087:

Fuji apples cost $3.00 per pound, and Golden Delicious apples cost $2.00 per pound. A childcare center purchases 30 pounds of a combination of the two types of apples for a total of $80. Which value could replace x in the table? 30 – g + g 3(30 – g) 30 – g(3) 30 – g + 3

Astrid1:

Given Fuji apple cost $3 per pound Golden Apple cost $2 per pound Purchase 30 pound in $80 Let quantity of fuji Apple is f and golden Apple is g So Equation's f+g=30.............(1) 3f+2g=80...........(2) After solve this equation we get f=20 and g=10 Put value of f in equation (1) and multiple by 3 New equation is 3×20+3(g)=90 =90-60-3(g) =30-3(g) This 30-3(g) value can put in the table. Source: https://www.assignmentexpert.com/homework-answers/mathematics/algebra/question-158763